# DS-VTON: High-Quality Virtual Try-on via Disentangled Dual-Scale Generation

### v1

@@ Abstract
Despite recent progress, most existing virtual try-on methods still struggle to simultaneously address two core challenges: accurately aligning the garment image with the target human body, and preserving fine-grained garment textures and patterns. In this paper, we propose DS-VTON, a dual-scale virtual try-on framework that explicitly disentangles these objectives for more effective modeling. DS-VTON consists of two stages: the first stage generates a low-resolution try-on result to capture the semantic correspondence between garment and body, where reduced detail facilitates robust structural alignment. The second stage introduces a residual-guided diffusion process that reconstructs high-resolution outputs by refining the residual between the two scales, focusing on texture fidelity. In addition, our method adopts a fully mask-free generation paradigm, eliminating reliance on human parsing maps or segmentation masks. By leveraging the semantic priors embedded in pretrained diffusion models, this design more effectively preserves the person’s appearance and geometric consistency. Extensive experiments demonstrate that DS-VTON achieves state-of-the-art performance in both structural alignment and texture preservation across multiple standard virtual try-on benchmarks.
Figure 1: Comparison between existing methods (Kim et al., 2024; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025) and our proposed DS-VTON. (a) Existing approaches typically adopt a single-scale pipeline with masked inputs, limiting their ability to capture full-body semantics and garment structure. (b) In contrast, DS-VTON employs a fully mask-free, dual-scale framework: the low-resolution stage focuses on garment-body alignment, while the high-resolution stage refines texture details via residual-guided denoising.
@@ 1 Introduction
Given a garment image and a person image, the goal of virtual try-on is to synthesize a photorealistic image of the person wearing the specified garment (Han et al., 2018). As a key enabling technology for online fashion and e-commerce, virtual try-on has attracted increasing attention in recent years (Choi et al., 2021; Ge et al., 2021b; Gou et al., 2023; Lee et al., 2022; Morelli et al., 2022, 2023; Choi et al., 2024; Chong et al., 2025; Zhou et al., 2025). This task involves two fundamental challenges: (1) accurately fitting the garment onto the human body, and (2) preserving fine-grained garment textures. Existing methods fall into two main categories: Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) and Diffusion Models (Ho et al., 2020; Rombach et al., 2022). Early GAN-based approaches (Choi et al., 2021; Ge et al., 2021a; Lee et al., 2022; Xie et al., 2023) typically follow a two-stage pipeline: a warping module first aligns the garment with the target pose, followed by a generation module to synthesize the final image. While warping helps preserve garment appearance, the subsequent generation stage often leads to detail loss due to imperfect feature fusion.
@@
Recent diffusion-based methods (Kim et al., 2024; Zhu et al., 2023; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025) address both structural alignment and texture preservation within a unified denoising process. Early diffusion steps tend to capture global structure, while later steps refine texture details. This progressive generation order aligns naturally with the requirements of virtual try-on. However, relying on a single-stage diffusion process remains inherently limited. In practice, many existing approaches still struggle to simultaneously ensure accurate garment-body alignment and high-fidelity texture reconstruction. The unified framework may fail to fully disentangle structure from detail, leading to compromised visual quality in the final output.
To address these challenges, we propose a dual-scale framework that explicitly decouples structural alignment and texture refinement. Motivated by the natural progression from structure to detail in diffusion processes, our pipeline first performs coarse alignment and subsequently enhances garment details in a second refinement stage. Specifically, our framework consists of two sequential stages:
  • Low-resolution stage: the model generates a coarse try-on result by suppressing high-frequency content and focusing on accurate structural alignment.
  • High-resolution stage: we extend a residual-guided denoising strategy that predicts the residual between the high-resolution image and its low-resolution counterpart. This enables the model to focus specifically on texture restoration, building upon the structurally aligned base from the first stage.
Beside, traditional virtual try-on methods rely on human parsing masks for spatial guidance, our approach adopts a mask-free strategy that eliminates this dependency by leveraging the strong semantic priors embedded in pretrained diffusion models. This design avoids errors caused by inaccurate masks and preserves the full-body appearance and geometry, allowing the model to adaptively infer garment regions. As a result, our method improves the accuracy of body structure modeling and better retains the individuality of the person’s appearance, leading to more realistic and personalized try-on results. Extensive experiments on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) validate the effectiveness of our method, demonstrating state-of-the-art performance both qualitatively and quantitatively.
Figure 2: Overview of our proposed method. The overall pipeline consists of a two-scale generation process: the low-resolution stage produces a coarse try-on result, which is then refined by the high-resolution stage. Both stages share the same network architecture. See Section 3 for further details.
@@ 2 Related works
Earlier methods (Choi et al., 2021; Ge et al., 2021a; Lee et al., 2022; Xie et al., 2023), which are based on Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), typically decompose the virtual try-on task into two stages: (1) warping the garment to align with the human body shape, and (2) integrating the warped garment with the human image to generate the final result. For instance, ACGPN (Yang et al., 2020) employs a warping module based on Thin-Plate Spline (TPS) (Duchon, 1977) to deform the garment. PFAFN (Ge et al., 2021b) proposes a parser-free method that guides the garment warping process using learned appearance flows. VITON-HD (Choi et al., 2021) introduces a specialized normalization layer and generator design to better handle garment-body misalignment during synthesis. However, a key limitation of GAN-based approaches is their constrained capacity in capturing both precise spatial alignment and fine-grained garment details. As a result, these methods often rely heavily on the warping stage to encode garment appearance early in the pipeline. Yet, the subsequent integration stage frequently introduces artifacts or detail loss, degrading the overall realism of the try-on result.
@@
With the rapid advancement of diffusion models (Ho et al., 2020; Rombach et al., 2022), a series of powerful virtual try-on methods have been proposed (Kim et al., 2024; Zhu et al., 2023; Choi et al., 2024; Zhou et al., 2025; Sun et al., 2024). Early approaches like DCI-VTON (Gou et al., 2023) follow a traditional two-stage pipeline: first warping the garment to match the body, then applying a diffusion model to blend it with the person image. More recent methods (Morelli et al., 2023; Kim et al., 2024; Zhu et al., 2023; Choi et al., 2024) adopt a single diffusion process that directly synthesizes the try-on result, shifting the focus to more effective conditioning strategies. For example, LaDI-VTON (Morelli et al., 2023) employs textual inversion to encode garment identity, while IDM-VTON (Choi et al., 2024) introduces a GarmentNet module for structural and appearance guidance. Leffa (Zhou et al., 2025) further proposes a Leffa loss to guide attention weights during the final 500 denoising steps, enhancing fine-grained texture recovery. Complementary to these methods, FitDiT (Jiang et al., 2024) explores an alternative direction. Built on SD3 (Esser et al., 2024), which adopts a DiT-based architecture (Peebles and Xie, 2023), it uses an aggressive rectangular mask to impose stronger spatial constraints, effectively mitigating the alignment issues faced by earlier mask-based pipelines. Despite notable progress, current diffusion-based methods still suffer from visible shortcomings. These include garment fragmentation artifacts due to imprecise human segmentation, as well as inaccurate rendering of fine patterns such as flowers or text. Addressing these open challenges remains a key direction.
3 Methodologies
@@
Given a person image $\mathbf{x}_{p}\in\mathbb{R}^{H\times W\times 3}$ and a garment image $\mathbf{x}_{g}\in\mathbb{R}^{H\times W\times 3}$, the virtual try-on task aims to generate a realistic output $\mathbf{x}_{r}\in\mathbb{R}^{H\times W\times 3}$, where the person appears to be wearing the given garment. Our pipeline requires no auxiliary conditions such as segmentation masks or human parsing maps. The only additional variables appear in the low-resolution stage, where the inputs $\mathbf{x}_{p}$ and $\mathbf{x}_{g}$ are downsampled to $\tilde{\mathbf{x}}_{p},\tilde{\mathbf{x}}_{g}\in\mathbb{R}^{h\times w\times 3}$, with $h=H/\sigma$, $w=W/\sigma$, and $\sigma$ denoting the downsampling ratio. These are used to generate a low-resolution try-on result $\tilde{\mathbf{x}}_{r}\in\mathbb{R}^{h\times w\times 3}$. For notational simplicity, we assume that $\tilde{\mathbf{x}}_{r}$ has been upsampled to the original resolution when used in the high-resolution stage, and we do not introduce a separate symbol for it. Further details on the low-resolution process and the choice of $\sigma$ are provided in Subsection 3.3.
@@
Previous virtual try-on methods (Kim et al., 2024; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025; Jiang et al., 2024) commonly rely on external human parsers to obtain body segmentation masks. These masks are then used to explicitly isolate the garment region in the person image, serving as spatial guidance during the try-on process. While this strategy is reasonable in conventional pipelines, it becomes less compelling in the context of diffusion-based models like Stable Diffusion (Rombach et al., 2022), which are pretrained on large-scale data and inherently encode strong semantic priors related to human structure and clothing layout. Modern pretrained diffusion models already demonstrate robust understanding of body geometry and clothing regions without requiring explicit mask supervision. Motivated by this observation, we adopt a fully mask-free design. Unlike prior work (Ge et al., 2021b), we do not distill pseudo-masks from human parsers. Instead, during both training and inference, the model directly takes a garment image and a person image as input, without relying on any parsing or segmentation guidance. This design offers two major advantages. First, segmentation masks may contain errors, either omitting parts of the garment or incorrectly covering body regions, leading to unrealistic or inconsistent outputs. Second, to avoid garment omission, large masks are sometimes used, which can obscure body regions and impair the model’s ability to reconstruct accurate body shape. In contrast, our mask-free strategy allows the model to adaptively infer the true garment region while preserving full-body geometry. This leads to more faithful body reconstruction and benefits downstream stages, ultimately improving the visual quality of the final try-on result. Please refer to Subsection 4.2 for qualitative comparisons.
@@ 3.1 Overview
In the low-resolution stage, we first use a VAE (Kingma, 2013) to encode the downsampled garment image $\tilde{\mathbf{x}}_{g}$, which is then passed to the reference U-Net. The downsampled person image $\tilde{\mathbf{x}}_{p}$ is also VAE-encoded and concatenated along the feature channel with a Gaussian noise tensor of the same shape. This combined latent is fed into the denoising U-Net to generate a low-resolution result $\tilde{\mathbf{x}}_{r}\in\mathbb{R}^{h\times w\times 3}$.
In the high-resolution stage, $\mathbf{x}_{g}$ and $\mathbf{x}_{p}$ are used in the same manner as in the low-resolution stage: $\mathbf{x}_{g}$ is passed to the reference U-Net, while $\mathbf{x}_{p}$ is encoded and concatenated with the latent, which is then input to the denoising U-Net. The main difference is the initialization of the latent. Specifically, instead of initializing $\mathbf{x}_{T}$ as standard Gaussian noise, we define the noisy latent input as:
@@ 3.2 Network architecture
We adopt a dual U-Net architecture, where the reference U-Net encodes garment features and integrates them into the main denoising U-Net through self-attention layers. This structure has proven effective for maintaining garment fidelity and enhancing visual quality (Choi et al., 2024), and is retained in our architecture. Following the design philosophy of (Chong et al., 2025), we remove all cross-attention layers in the U-Net, relying solely on self-attention mechanisms. This simplification improves performance and efficiency, as confirmed by our ablation studies (refer to Appendix B.1). All garment-related information is delivered through the reference U-Net, eliminating the need for additional conditioning pathways. To further improve computational efficiency, we execute the reference U-Net only once per sample, treating it purely as a conditioning module, in line with (Li et al., 2024).
We initialize our U-Net weights using those of Stable Diffusion 1.5 (Rombach et al., 2022). While SDXL (Podell et al., 2025) offers more powerful generative capabilities, the goal of our work is to propose a lightweight yet effective framework. Hence, we retain SD1.5 as our backbone. We also conducted preliminary experiments based on transformer architectures such as SD3 and SD3.5, with additional discussion provided in Appendix C. In conclusion, SD1.5 offers an optimal trade-off between model simplicity and try-on performance, and forms the foundation of our pipeline.
@@ 3.3 Low-resolution stage
Low-resolution images inherently emphasize structural information due to the suppression of fine-grained textures. Thus, we first introduce a low-resolution stage to guide the overall generation process. This stage serves two main purposes. First, it enables the model to capture the human body shape and garment category, ensuring accurate structural alignment between the person and the clothing. Second, it provides coarse but semantically reliable garment structure, such as the placement of stripes or patterns, offering guidance for fine-detail reconstruction in the high-resolution stage.
In this stage, we first downsample the garment image $\mathbf{x}_{g}$ and the person image $\mathbf{x}_{p}$ by the downsampling ratio $\sigma$, and then perform denoising from pure Gaussian noise $\epsilon$ to generate the low-resolution result $\tilde{\mathbf{x}}_{r}$. Both training and inference follow the standard diffusion process used in Stable Diffusion (Rombach et al., 2022; Ho et al., 2020; Song et al., 2020).
The only hyperparameter in this stage is the downsampling ratio $\sigma$. We consider three values: $\sigma=1,2,$ and $4$. When $\sigma=1$, the low-resolution stage operates at the same resolution as the high-resolution stage; $\sigma=2$ corresponds to downsampling from $768\times 1024$ to $384\times 512$, and so on. Based on our observations, downsampling ratios of $\sigma=2$ and $\sigma=4$ both enhance human body structure understanding, while $\sigma=2$ offers the most reliable structural guidance for accurate garment reconstruction. Quantitative and qualitative comparisons of different $\sigma$ values are presented in Subsection 4.3, and $\sigma=2$ is used in all experiments throughout this work.

### v2

# DS-VTON: An Enhanced Dual-Scale Coarse-to-Fine Framework for Virtual Try-On

@@ Abstract
Despite recent progress, most existing virtual try-on methods still struggle to simultaneously address two core challenges: accurately aligning the garment image with the target human body, and preserving fine-grained garment textures and patterns. These two requirements map directly onto a coarse-to-fine generation paradigm, where the coarse stage handles structural alignment and the fine stage recovers rich garment details. Motivated by this observation, we propose DS-VTON, an enhanced dual-scale coarse-to-fine framework that tackles the try-on problem more effectively. DS-VTON consists of two stages: the first stage generates a low-resolution try-on result to capture the semantic correspondence between garment and body, where reduced detail facilitates robust structural alignment. In the second stage, a blend-refine diffusion process reconstructs high-resolution outputs by refining the residual between scales through noise–image blending, emphasizing texture fidelity and effectively correcting fine-detail errors from the low-resolution stage. In addition, our method adopts a fully mask-free generation strategy, eliminating reliance on human parsing maps or segmentation masks. Extensive experiments show that DS-VTON not only achieves state-of-the-art performance but consistently and significantly surpasses prior methods in both structural alignment and texture fidelity across multiple standard virtual try-on benchmarks. More visualizations are available in our project page.
Figure 1: (a) DS-VTON results across diverse scenarios. (b) Existing methods (Kim et al., 2024; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025) adopt a single-scale pipeline with masked inputs, limiting their ability to capture full-body semantics and garment structure. (c) In contrast, DS-VTON adopts an enhanced dual-scale coarse-to-fine framework combined with a mask-free strategy.
@@ 1 Introduction
Given a garment image and a person image, the goal of virtual try-on is to synthesize a photorealistic image of the person wearing the specified garment (Han et al., 2018). As a key enabling technology for online fashion and e-commerce, virtual try-on has attracted increasing attention in recent years (Choi et al., 2021; Ge et al., 2021b; Gou et al., 2023; Lee et al., 2022; Zhang et al., 2024; Morelli et al., 2022, 2023; Choi et al., 2024; Chong et al., 2025; Zhou et al., 2025).
@@
This task involves two fundamental challenges: (1) accurately fitting the garment onto the human body, and (2) preserving fine-grained garment textures. Existing methods fall into two main categories: Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) and Diffusion Models (Ho et al., 2020; Rombach et al., 2022). Early GAN-based approaches (Choi et al., 2021; Ge et al., 2021a; Lee et al., 2022; Xie et al., 2023) typically follow a two-stage pipeline: a warping module first aligns the garment with the target pose, followed by a generation module to synthesize the final image. While warping helps preserve garment appearance, the subsequent generation stage often leads to detail loss due to imperfect feature fusion.
@@
Recent diffusion-based methods (Kim et al., 2024; Zhu et al., 2023; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025) have become popular. The denoising process naturally progresses from coarse to fine: early steps capture global structure, while later steps refine texture details (Balaji et al., 2022; Choi et al., 2022). This progressive generation order aligns well with the requirements of virtual try-on, where both alignment and texture fidelity are essential. However, relying on a single-stage diffusion process remains inherently limited. In practice, existing approaches still struggle to ensure accurate garment-body alignment and high-fidelity detail reconstruction. Without explicitly disentangling structure from detail, the unified framework often produces compromised visual quality.
To overcome the limitations of conventional single-stage denoising, we introduce a dual-scale coarse-to-fine framework that explicitly separates global structure alignment from fine-grained texture restoration. Low-resolution stage: In the first stage, the model generates a coarse try-on result by suppressing high-frequency content, emphasizing structural alignment through the low-resolution representation. High-resolution stage: In the second stage, a blend-refine diffusion process explicitly transforms the low-resolution output into high resolution, restoring fine textures and correcting fine-detail errors from the first stage.
Beside, traditional virtual try-on methods rely on human parsing masks for spatial guidance, our approach adopts a fully mask-free strategy that eliminates this dependency by leveraging the strong semantic priors embedded in pretrained diffusion models. Our main contributions are as follows:
  • We propose a novel dual-scale, mask-free framework that enhances the coarse-to-fine process and is particularly well-suited for the try-on task.
  • We introduce a blend-refine diffusion process that explicitly bridges two complex distributions, enabling controllable transition from coarse alignment to fine-detail restoration.
  • Extensive experiments on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) demonstrate state-of-the-art performance, validating the effectiveness of our method both qualitatively and quantitatively.
@@ 2 Related works
Earlier methods (Choi et al., 2021; Ge et al., 2021a; Lee et al., 2022; Xie et al., 2023), which are based on Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), typically decompose the virtual try-on task into two stages: (1) warping the garment to align with the human body shape, and (2) integrating the warped garment with the human image to generate the final result. For instance, ACGPN (Yang et al., 2020) employs a warping module based on Thin-Plate Spline (TPS) (Duchon, 1977) to deform the garment. PFAFN (Ge et al., 2021b) proposes a parser-free method that guides the garment warping process using learned appearance flows. VITON-HD (Choi et al., 2021) introduces a specialized normalization layer and generator design to better handle garment-body misalignment during synthesis.
However, a key limitation of GAN-based approaches is their constrained capacity in capturing both precise spatial alignment and fine-grained garment details. As a result, these methods often rely heavily on the warping stage to encode garment appearance early in the pipeline. Yet, the subsequent integration stage frequently introduces artifacts or detail loss, degrading the overall realism of the try-on result.
@@
With the rapid advancement of diffusion models (Ho et al., 2020; Rombach et al., 2022), powerful virtual try-on methods have emerged (Kim et al., 2024; Zhu et al., 2023; Choi et al., 2024; Zhou et al., 2025; Sun et al., 2024). Early methods like DCI-VTON (Gou et al., 2023) use a two-stage pipeline: warping the garment to match the body, then blending it with the person image using a diffusion model. DT-VTON (Zhang et al., 2024) splits the task into structural alignment and texture replacement. More recent methods (Morelli et al., 2023; Kim et al., 2024; Zhu et al., 2023; Choi et al., 2024) employ a single diffusion process for direct try-on synthesis, with improved conditioning strategies. For instance, LaDI-VTON (Morelli et al., 2023) uses textual inversion to encode garment identity, while IDM-VTON (Choi et al., 2024) introduces GarmentNet for structural and appearance guidance. Leffa (Zhou et al., 2025) proposes a Leffa loss to guide attention weights during the final 500 denoising steps for better texture recovery. Complementary to these, FitDiT (Jiang et al., 2024) adopts a DiT-based architecture (Peebles and Xie, 2023) with an aggressive rectangular mask to address alignment issues in mask-based pipelines.
Despite progress, diffusion-based methods still face challenges, such as garment fragmentation due to imprecise segmentation and poor rendering of fine patterns like flowers or text, which remain key areas for improvement.
Figure 2: Upper panel: Two-scale generation pipeline. A low-resolution stage produces a coarse try-on result, then refined by a high-resolution stage; both stages share the same network architecture (see Section 3). Lower panel: Results with different settings; ours uses $\sigma=2$ and $\alpha=\beta=\tfrac{1}{2}$(see Subsections 3.3 and 3.4). With proper two-stage settings, the second stage leverages the reliable coarse structure from the first stage to correct fine-detail errors and generate high-quality try-on results.
3 Methods
@@
Given a person image $\mathbf{x}_{p}\in\mathbb{R}^{H\times W\times 3}$ and a garment image $\mathbf{x}_{g}\in\mathbb{R}^{H\times W\times 3}$, the virtual try-on task generates a realistic output $\mathbf{x}_{r}\in\mathbb{R}^{H\times W\times 3}$, where the person wears the given garment. In the low-resolution stage, the inputs $\mathbf{x}_{p}$ and $\mathbf{x}_{g}$ are downsampled to $\tilde{\mathbf{x}}_{p},\tilde{\mathbf{x}}_{g}\in\mathbb{R}^{h\times w\times 3}$, with $h=H/\sigma$, $w=W/\sigma$, and $\sigma$ denoting the downsampling ratio. These are used to generate a low-resolution result $\tilde{\mathbf{x}}_{r}\in\mathbb{R}^{h\times w\times 3}$. For simplicity, we assume that $\tilde{\mathbf{x}}_{r}$ is upsampled to the original resolution in the high-resolution stage, and do not introduce a separate symbol.
@@
Previous virtual try-on methods (Kim et al., 2024; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025; Jiang et al., 2024) typically use external human parsers to generate body segmentation masks for garment localization. However, in diffusion-based models like Stable Diffusion (Rombach et al., 2022), which already encode strong human-structure priors, this is unnecessary. We therefore adopt a fully mask-free design: the model directly consumes the garment and person images during training and inference, without parser-based pseudo-masks or segmentation guidance.
@@ 3.1 Overview
In the low-resolution stage, images inherently emphasize structural information because fine-grained textures are suppressed. This stage therefore serves two main purposes. First, it enables the model to capture the human body shape and garment category, ensuring accurate structural alignment between the person and the clothing. Second, it provides coarse but semantically reliable garment structure, such as the placement of stripes or patterns, offering guidance for fine-detail reconstruction in the high-resolution stage.
In the high-resolution stage, the key challenge is how to effectively utilize the low-resolution result $\tilde{\mathbf{x}}_{r}$. Here, we introduce our blend-refine diffusion process, which defines the noisy latent input as
@@ 3.2 Network architecture
We adopt a dual U-Net architecture, where the reference U-Net encodes garment features and integrates them into the main denoising U-Net via self-attention layers, a structure proven effective for garment fidelity and visual quality (Choi et al., 2024). Following the approach of (Chong et al., 2025), we remove all cross-attention layers, relying solely on self-attention, which improves performance and efficiency, as shown by our ablation studies (see Appendix A.2.1). To enhance computational efficiency, we execute the reference U-Net once per sample, using it solely as a conditioning module, as in (Li et al., 2024).
We initialize our U-Net weights using those from Stable Diffusion 1.5 (Rombach et al., 2022). While SDXL (Podell et al., 2025) offers greater generative power, our goal is a lightweight yet effective framework, so we retain SD1.5 as the backbone. Preliminary experiments with transformer-based architectures like SD3 and SD3.5 are discussed in Appendix A.3. In conclusion, SD1.5 strikes the optimal balance between model simplicity and try-on performance, forming the foundation of our pipeline.
@@ 3.3 Low-resolution stage
In this stage, we first downsample the garment image $\mathbf{x}_{g}$ and the person image $\mathbf{x}_{p}$ by the downsampling ratio $\sigma$. The downsampled garment image $\tilde{\mathbf{x}}_{g}$ is encoded by a VAE (Kingma, 2013) and then fed into the reference U-Net. The downsampled person image $\tilde{\mathbf{x}}_{p}$ is also VAE-encoded and concatenated along the feature channels with a Gaussian noise tensor of the same shape. This combined latent is passed to the denoising U-Net to generate a low-resolution result $\tilde{\mathbf{x}}_{r}\in\mathbb{R}^{h\times w\times 3}$. Both training and inference follow the standard diffusion process used in Stable Diffusion (Rombach et al., 2022; Ho et al., 2020; Song et al., 2020).
The only hyperparameter in this stage is the downsampling ratio $\sigma$, with $\sigma\in\{1,2,4\}$. As illustrated in Figure 2, when $\sigma=1$, the low-resolution stage operates at the same resolution as the high-resolution stage, violating the purpose of leveraging lower resolution for improved structural modeling and introducing artifacts due to difficulty. $\sigma=2$ corresponds to downsampling from $768\times 1024$ to $384\times 512$, and so on. Our observations show that both $\sigma=2$ and $\sigma=4$ enhance human-body structural understanding, but, as also evident in Figure 2, $\sigma=4$ sacrifices structural detail, making $\sigma=2$ the most reliable choice for accurate garment reconstruction. Quantitative and qualitative comparisons across different $\sigma$ values are provided in Subsection 4.3, and $\sigma=2$ is adopted for all experiments in this work.